A new vaccine was tested to see if it could prevent the ear infections that many infants suffer from. Babies about a year old were randomly divided into two groups. One group received​ vaccinations, and the other did not. The following​ year, only 329 of 2456 vaccinated children had ear​ infections, compared to 496 of 2448 unvaccinated children. Complete parts​ a) through​ c) below.

Let p1 be the sample proportion of success in the unvaccinated​ group, and let p2 be the sample proportion of success in the vaccinated group. Find the​ 95% confidence interval for the difference in rates of ear​ infection, p1−p2.

The confidence interval is (___%,___%).

there is a 95% confidence, based on these sample, that the proportion of American woman, age 65and older who suffer from arthritis is between ____% and ___% more or less than the proportion of American men of the same age who suffer
Answer #1

Sample proportion Sample Sample number Size of events Sample 1 n1 = 2448 X1 =496 P1 = X1/n1 =(496/2448) = 0.2026 Sample 2 n2

Ans :

The confidence interval is (4.77 %, 8.95 %).

there is a 95% confidence, based on these sample, that the proportion of American woman, age 65and older who suffer from arthritis is between 4.77% and 8.95% more or less than the proportion of American men of the same age who suffer